# Faster Reinforcement Learning by Freezing Slow States

### v3

@@ Introduction
Markov decision processes (MDPs) offer a foundational framework for modeling sequential decision making under uncertainty. In many real-world applications, decisions must be made at high frequencies over extended time horizons. From a modeling standpoint, this translates into MDPs with discount factors close to one (which capturing the long-term nature of the objective). However, such long-horizon problems can pose computational challenges: standard algorithms like value iteration (VI), which is a core building block in many reinforcement learning (RL) algorithms, rely on the contraction property of the Bellman operator to guarantee convergence. As the discount factor approaches one, this contraction weakens, leading to slow convergence and degraded performance within a given computational budget (Bertsekas and Tsitsiklis, 1996; Jiang et al., 2015).
@@
Solving MDPs that fully capture both fast and slow dynamics at the natural decision frequency typically requires long effective horizons, which can pose serious computational burdens. To manage complexity, practitioners often resort to simplifying assumptions, such as fixing or omitting slow states altogether. While this reduces computational burden, it can degrade policy quality by discarding important information. This paper introduces and formalizes the notion of fast-slow MDPs, and studies the computational implications of a novel multi-scale approximation strategy called frozen-state value iteration (FSVI).
@@
    We propose a frozen-state approximation to the reformulated MDP, along with an associated frozen-state value iteration (FSVI) algorithm, where the slow state is frozen in the lower-level problem, while each period in the upper-level problem “releases” the slow state. Computational benefits arise in several ways: (1) frozen states simplify the dynamics of the lower-level MDP (fewer successor states), (2) the discount factor in the upper-level problem is more favorable due to the reuse of the lower-level policy, which is computed only once, (3) the lower-level MDP thus becomes separable into independent MDPs, opening the door to speedups via parallel computation. Solving the frozen-state approximation gives a policy that switches between one action from upper-level policy and $T-1$ actions from the lower-level policy. We give a theoretical analysis that upper bounds the expected regret from applying this policy compared to the optimal policy.
@@
The options (and subgoals) framework is largely motivated by robotics and navigation-related tasks, while we are particularly interested in solving problems that arise in the operations research and operations management domains. The problems that we study do not decompose naturally into “subgoals” — leading us to identify and focus on the fast-slow structure, which does indeed arise naturally for many problems of interest.
Another work that is related to the options framework is Song and Xu (2020), who divide finite-horizon MDPs into two sub-problems along the time horizon, and concatenate their optimal solutions to generate an overall solution. Our paper also, in a sense, divides MDPs along the time horizon, but our work is quite different from that of Song and Xu (2020) in that we work on infinite horizon problems and convert them into auxiliary problems that operate on a slower timescale, which takes advantage of reusable lower-level policies. More importantly, the various methods we propose all build on the idea of freezing certain states to reduce computational cost, which is unique to our approach and to our knowledge, this is a novel direction that has not been proposed before.
Finally, our work contributes to the broader literature on leveraging MDP structure to improve reinforcement learning and approximate dynamic programming algorithm. Prior research has exploited structural properties of the value function, such as convexity (Pereira and Pinto, 1991; Philpott and Guan, 2008; Nascimento and Powell, 2009; Benjaafar et al., 2022) and monotonicity (Papadaki and Powell, 2002; Jiang and Powell, 2015a), as well as structural features of the optimal policy (Kunnumkal and Topaloglu, 2008). Additional examples include factored MDPs (Osband and Van Roy, 2014) and weakly-coupled systems (Killian et al., 2021; El Shar and Jiang, 2023; Brown and Smith, 2025; Nadarajah and Cire, 2025). Our contribution to this line of work lies in identifying a new form of MDP structure, which then informs the design of a new algorithmic approach.
@@
$(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$ Output policy of FSVI after $k$ iterations
@@
Denote $(\mu,\bm{\pi})$ a $T$-horizon policy, which is a sequence of $T$ policies $(\mu,\pi_{1},\ldots,\pi_{T-1})$, $\mu:\mathcal{S}\rightarrow\mathcal{A}$, $\pi_{t}:\mathcal{S}\rightarrow\mathcal{A}$ and $\bm{\pi}=(\pi_{1},\ldots,\pi_{T-1})$. Following $(\mu,\bm{\pi})$ means that we take the first action according to $\mu$ and then next $T-1$ actions according to $\bm{\pi}$. Given any state $s_{0}$, the $T$-period reward function (of the base model) associated with $(\mu,\bm{\pi})$ is written as:
@@
We view the problem from period $1$ to period $T$ as the “lower level” of the frozen-state approximation. [^4] This corresponds to the periods relevant to $\bm{\pi}$ from $(\mu,\bm{\pi})$ in the hierarchical reformulation (8), whose structure the frozen-state approximation mimics. To form the lower-level problem of the frozen-state approximation, we consider this $T-1$ period problem in isolation:
@@
Note that using $J_{T}^{*}\equiv 0$ is aligned with the fact that the lower-level problem (9) contains $T-1$ periods. It thus follows that $J_{1}^{\tilde{\bm{\pi}}^{*}}=J_{1}^{*}$, meaning that we can use $J_{1}^{*}$ to represent the value of the lower-level problem. We will see in the next section that the upper-level problem depends on an approximation of $J_{1}^{\tilde{\bm{\pi}}^{*}}$, which is the value of the lower-level policy in $T-1$ periods (and zero terminal value after that).
@@
Recall that the optimal lower-level policy (for the frozen-state model) is denoted $\tilde{\bm{\pi}}^{*}$ and its optimal value is $J_{1}^{*}$. Let $\tilde{\mu}^{*}$ be the optimal upper-level policy corresponding to these inputs, i.e., the policy that is greedy with respect to $V^{*}(s_{0},J_{1}^{*},\tilde{\bm{\pi}}^{*})$. Thus, $(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})$ is the resulting $T$-periodic policy from the frozen-state hierarchical approximation; we refer to it as the $T$-periodic frozen-state policy.
@@
    Solve the lower-level MDP with frozen states to obtain a policy $\tilde{\bm{\pi}}^{*}$ and its value $J_{1}^{*}$. Since the lower-level problem is a finite horizon MDP, it can be solved exactly using $T-1$ steps of VI. An alternative way of viewing the lower-level problem is that it is $|\mathcal{X}|$ independent MDPs, which we can potentially solve in parallel. Note that Algorithm 2 emphasizes this by looping over the slow state $x\in\mathcal{X}$ in Line 1.
@@
We denote the resulting value function approximation after $k$ iterations of value iteration as $V^{k}$, from which we obtain a policy $\tilde{\mu}^{k}$. The $T$-periodic policy output by FSVI is formed by combining $\tilde{\mu}^{k}$ with the optimal finite-horizon policy $\tilde{\bm{\pi}}^{*}$ from the lower-level MDP: $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$.
@@
1 for $k^{\prime}=1,2,\ldots,k$ do
2       for $s$ in the state space $\mathcal{S}$ do
3             $U^{k^{\prime}}(s)=\max_{a}r(s,a)+\gamma\,\mathbb{E}\bigl{[}U^{k^{\prime}-1}(f(s,a,w))\bigr{]}.$
@@
7 for $s$ in the state space $\mathcal{S}$ do
8       $\nu^{k}(s)=\operatorname*{arg\,max}_{a}r(s,a)+\gamma\,\mathbb{E}\bigl{[}U^{k}(f(s,a,w))\bigr{]}$.
@@
Output: Approximation of the $T$-periodic frozen-state policy $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$ and $J_{1}^{*}$.
@@
2for each slow state $x\in\mathcal{X}$ do
3       for $t=T-1,T-2,\ldots,1$ do
4             for each fast state $y\in\mathcal{Y}$ do
5                   $J^{*}_{t}(x,y)=\max_{a}r(x,y,a)+\gamma\,\mathbb{E}\bigl{[}J^{*}_{t+1}(x,f_{\mathcal{Y}}(x,y,a,w))\bigr{]}$.
6                   $\tilde{\pi}_{t}^{*}(x,y)=\operatorname*{arg\,max}_{a}r(x,y,a)+\gamma\,\mathbb{E}\bigl{[}J^{*}_{t+1}(x,f_{\mathcal{Y}}(x,y,a,w))\bigr{]}$.
@@
12for $k^{\prime}=1,2,\ldots,k$ do
13       for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
14             $V^{k^{\prime}}(x_{0},y_{0},J^{*}_{1},\tilde{\bm{\pi}}^{*})=\max_{a}\mathbb{E}\bigl{[}\tilde{R}(s_{0},a,J^{*}_{1})+\gamma^{T}V^{i-1}(x_{T},y_{T},J^{*}_{1},\tilde{\bm{\pi}}^{*})\bigr{]}$.
@@
18 for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
19       $\tilde{\mu}^{k}(x_{0},y_{0})=\operatorname*{arg\,max}_{a}\mathbb{E}\bigl{[}\tilde{R}(s_{0},a,J^{*}_{1})+\gamma^{T}V^{k}(x_{T},y_{T},J^{*}_{1},\tilde{\bm{\pi}}^{*})\bigr{]}$.
@@
$\text{Cost}_{\text{FSVI}}(k)=\;\mathcal{O}\!\Bigl{(}\,\underbrace{|\mathcal{X}|\,|\mathcal{Y}|^{2}\,|\mathcal{A}|\,T\;+\;|\mathcal{S}|^{2}\,T}_{\begin{subarray}{c}\text{one-time lower-level}\\ \text{pre\textendash processing cost}\end{subarray}}\;+\;\underbrace{k\,|\mathcal{S}|^{2}\,|\mathcal{A}|}_{\begin{subarray}{c}\text{upper-level}\\ \text{iterations}\end{subarray}}\,\Bigr{)}.$
@@
[…]
where $\bm{0}$ is the all-zero value function. The difference between (16) and (17) can be interpreted as follows: in the former, we follow a lower-level policy that is aware of a terminal value $U^{*}$ (but exclude that value when defining the $T$-horizon reward), while in the latter, we follow a lower-level policy that sees zero terminal reward at the end of the $T-1$ periods.
@@
As a follow-up comment to Remark 2, notice that $V^{*}(s_{0},J_{1}^{*},\tilde{\bm{\pi}}^{*})$ does not directly enter the regret definition, as $V^{*}(s_{0},J_{1}^{*},\tilde{\bm{\pi}}^{*})$ is just the optimal value of the frozen-state approximation, not the value of its implied greedy policy $\tilde{\mu}^{*}$ when evaluated in the base model. However, the regret of course depends on $V^{*}(s_{0},J_{1}^{*},\tilde{\bm{\pi}}^{*})$ indirectly, because $\tilde{\mu}^{*}$ depends on $V^{*}(s_{0},J_{1}^{*},\tilde{\bm{\pi}}^{*})$.
@@ A General Lemma
[…]
@@
Let $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$ be the resulting $T$-periodic policy after running FSVI for $k$ iterations. The regret incurred when running $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$ in the base model satisfies
$\mathcal{R}(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})\leq$ $\biggl{(}\frac{2\gamma^{T}}{(1-\gamma^{T})^{2}}+\frac{2}{1-\gamma^{T}}\biggr{)}\epsilon_{r}(\gamma,\alpha,d_{\mathcal{Y}},\mathbf{L},T)$
@@
The above procedure can be converted into a simple visualization. In Figure 5, we plot regret versus computational cost. This is done for both FSVI for different values of $T$ and also for standard value iteration (Base VI). Each colored curve traces the regret-cost trade-off generated by varying $k$ for a distinct $T$ (though $k$ is not shown). Imposing the computational budget corresponds to drawing the vertical line $x=B$ and where this line intersects a curve, one can read off the best attainable regret given budget $B$ for the algorithm represented by that curve. By scanning the intersection points and selecting the one with the lowest regret, one can identify the best algorithm to use, whether it is FSVI with a particular $T$ or simply Base VI. Hence, the plot allows us to visually solve the constrained optimization problem of minimizing regret under a given computational budget. This exercise leads us to interesting takeaways illustrated in Figure 5:
@@
Input: State distribution $\rho$, number of iterations $k$, number of state samples $N$, number of next-state samples $M$, and a regression procedure Regress.
@@
3       Sample a dataset $\mathcal{D}=\{s^{(i)}\}_{i=1}^{N}\sim\rho$
4       for $s^{(i)}\in\mathcal{D}$ do
5             $u^{(i)}=\max_{a}\,r(s^{(i)},a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k^{\prime}-1}\bigl{(}s^{(i,a,j)}\bigr{)}$, where $s^{(i,a,j)}$ is sampled from the generative model, conditional on state $s^{(i)}$ and action $a$.
@@ 3       Sample a dataset $\mathcal{D}=\{s^{(i)}\}_{i=1}^{N}\sim\rho$
7       $U^{k^{\prime}}=\texttt{Regress}(\{(s^{(i)},u^{(i)})\}_{i=1}^{N})$
@@
9 for $s\in\mathcal{S}$ do
10       $\nu^{k}(s)=\operatorname*{arg\,max}_{a}\,r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k}\bigl{(}s^{(a,j)}\bigr{)}$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
@@
Input: State distribution $\rho$, number of iterations $k$, number of state samples $N$, number of next-state samples $M_{l}$ (lower level), $M_{u}$ (upper level), and a regression procedure Regress.
Output: Approximation of the $T$-periodic frozen-state policy $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$ and $\tilde{J}_{1}$.
1 Initialize $\tilde{J}_{T}\equiv 0$.
@@
3for $t=T-1,T-2,\ldots,1$ do
4       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
5       for each $(x^{(i)},y^{(i)})\in\mathcal{D}$ do
6             $J_{t}^{(i)}=\max_{a\in\mathcal{A}}\,r\bigl{(}x^{(i)},y^{(i)},a\bigr{)}+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}\tilde{J}_{t+1}\bigl{(}x^{(i)},y^{(i,a,j)}\bigr{)}$ where $y^{(i,a,j)}$ is sampled from the generative model, conditional on state $(x^{(i)},y^{(i)})$ and action $a$.
@@ 4       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
8       $\tilde{J}_{t}=\texttt{Regress}\Bigl{(}\bigl{\{}\bigl{(}(x^{(i)},y^{(i)}),J_{t}^{(i)}\bigr{)}\bigr{\}}_{i=1}^{N}\Bigr{)}$
@@
10Let $\tilde{\bm{\pi}}^{*}=(\pi_{1}^{*},\ldots\pi^{*}_{T-1})$, with $\pi_{t}^{*}(s)=\operatorname*{arg\,max}_{a}\,r(s,a)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}\,\tilde{J}_{t+1}\bigl{(}s^{(a,j)}\bigr{)}$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$. Initialize $\hat{V}^{0}$ arbitrarily. for $k^{\prime}=1,2,\ldots,k$ do
11       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
12       for each $(x^{(i)},y^{(i)})\in\mathcal{D}$ do
13             ${V}^{(k^{\prime})}=\max_{a\in\mathcal{A}}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}{V}^{k^{\prime}-1}\left(x^{(i,a,j)},y^{(i,a,j)}\right)\right]$ where $x_{T}^{(i,a,j)}$, $y_{T}^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled using a $T$-step transition conditioned on $x^{(i)}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $\tilde{J}_{1}$.
@@ 11       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
15       $\tilde{V}^{k^{\prime}}=\texttt{Regress}\Bigl{(}\bigl{\{}\bigl{(}(x^{(i)},y^{(i)}),{V}^{(i)}\bigr{)}\bigr{\}}_{i=1}^{N}\Bigr{)}$
@@
17for $s_{0}=(x_{0},y_{0})\in\mathcal{S}$ do
18       $\tilde{\mu}^{k}(x_{0},y_{0})=\operatorname*{arg\,max}_{a\in\mathcal{A}}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}\tilde{V}^{k}\left(x_{T}^{(i,a,j)},y_{T}^{(i,a,j)}\right)\right]$ where $x_{T}^{(i,a,j)}$, $y_{T}^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled from a $T$-step transition as above.
@@
  • {Lin, Deep}-FVI: Fitted value iteration (Munos and Szepesvári, 2008) using either linear function approximator (Lin-FVI) or a three-layer neural network function approximator (Deep-FVI). This algorithm is also known as least-square Monte Carlo, popularized by Longstaff and Schwartz (2001) and frequently used in the operations literature (see, e.g., Nadarajah et al. (2017)).
@@
We test these algorithms across three domains, summarized in Table 2. Each domain uses a high discount factor of $\gamma=0.995$, corresponding to an effective planning horizon of approximately $2001/(1-\gamma)=$ 200 steps. We describe each experiment below; in each section, we use self-contained notation. Detailed experimental settings are given in Appendix B.

### v4

@@ Introduction
Markov decision processes (MDPs) offer a foundational framework for modeling sequential decision making under uncertainty. In many real-world applications, decisions must be made at high frequencies over extended time horizons. From a modeling standpoint, this translates into MDPs with discount factors close to one (which captures the long-term nature of the objective). However, such long-horizon problems can pose computational challenges: standard algorithms like value iteration (VI), which is a core building block in many reinforcement learning (RL) algorithms, rely on the contraction property of the Bellman operator to guarantee convergence. As the discount factor approaches one, this contraction weakens, leading to slow convergence and degraded performance within a given computational budget (Bertsekas and Tsitsiklis, 1996; Jiang et al., 2015).
@@
Solving MDPs that fully capture both fast and slow dynamics at the natural decision frequency typically requires long effective horizons, which can pose serious computational burdens. To manage complexity, practitioners often resort to simplifying assumptions, such as fixing or omitting slow states altogether. While this reduces computational burden, it can degrade policy quality by discarding important information. This paper introduces and formalizes the notion of fast-slow MDPs, and studies the computational implications of a novel approximation strategy called frozen-state value iteration (FSVI).
@@
    We propose a frozen-state approximation to the reformulated MDP, along with an associated frozen-state value iteration (FSVI) algorithm, where the slow state is frozen in the lower-level problem, while each period in the upper-level problem “releases” the slow state. Computational benefits arise in several ways: (1) frozen states simplify the dynamics of the lower-level MDP (fewer successor states), (2) the discount factor in the upper-level problem is more favorable due to the reuse of the lower-level policy, which is computed only once, (3) the lower-level MDP thus becomes separable into independent MDPs, opening the door to speedups via parallel computation. Solving the frozen-state approximation gives a policy that switches between one action from upper-level policy and T1T-1 actions from the lower-level policy. We give a theoretical analysis that upper bounds the expected regret from applying this policy compared to the optimal policy.
@@
The options (and subgoals) framework is largely motivated by robotics and navigation-related tasks, while we are particularly interested in solving problems that arise in the operations research and operations management domains. The problems that we study do not decompose naturally into “subgoals,” leading us to identify and focus on the fast-slow structure, which does indeed arise naturally for many problems of interest.
Another work that is related to the options framework is Song and Xu (2020), who divide finite-horizon MDPs into two sub-problems along the time horizon, and concatenate their optimal solutions to generate an overall solution. Our paper also, in a sense, divides MDPs along the time horizon, but our work is quite different from that of Song and Xu (2020) in that we work on infinite horizon problems and convert them into auxiliary problems that operate on a slower timescale, which takes advantage of reusable lower-level policies. More importantly, the various methods we propose all build on the idea of freezing certain states to reduce computational cost, which is unique to our approach and, to our knowledge, this is a novel direction that has not been proposed before.
Finally, our work contributes to the broader literature on leveraging MDP structure to improve reinforcement learning and approximate dynamic programming algorithms. Prior research has exploited structural properties of the value function, such as convexity (Pereira and Pinto, 1991; Philpott and Guan, 2008; Nascimento and Powell, 2009; Benjaafar et al., 2022) and monotonicity (Papadaki and Powell, 2002; Jiang and Powell, 2015a), as well as structural features of the optimal policy (Kunnumkal and Topaloglu, 2008). Additional examples include factored MDPs (Osband and Van Roy, 2014) and weakly-coupled systems (Killian et al., 2021; El Shar and Jiang, 2023; Brown and Smith, 2025; Nadarajah and Cire, 2025). Our contribution to this line of work lies in identifying a new form of MDP structure, which then informs the design of a new algorithmic approach.
@@
[…]
@@
Denote $(\mu,\bm{\pi})$ a $T$-horizon policy, which is a sequence of $T$ policies $(\mu,\pi_{1},\ldots,\pi_{T-1})$, $\mu:\mathcal{S}\rightarrow\mathcal{A}$, $\pi_{t}:\mathcal{S}\rightarrow\mathcal{A}$ and $\bm{\pi}=(\pi_{1},\ldots,\pi_{T-1})$. Following $(\mu,\bm{\pi})$ means that we take the first action according to $\mu$ and then next T1T-1 actions according to $\bm{\pi}$. Given any state $s_{0}$, the $T$-period reward function (of the base model) associated with $(\mu,\bm{\pi})$ is written as:
@@
We view the problem from period $1$ to period $T$ as the “lower level” of the frozen-state approximation. [^4] This corresponds to the periods relevant to $\bm{\pi}$ from $(\mu,\bm{\pi})$ in the hierarchical reformulation (8), whose structure the frozen-state approximation mimics. To form the lower-level problem of the frozen-state approximation, we consider this T1T-1 period problem in isolation:
@@
Note that using $J_{T}^{*}\equiv 0$ is aligned with the fact that the lower-level problem (9) contains T1T-1 periods. It thus follows that $J_{1}^{\tilde{\bm{\pi}}^{*}}=J_{1}^{*}$, meaning that we can use $J_{1}^{*}$ to represent the value of the lower-level problem. We will see in the next section that the upper-level problem depends on an approximation of $J_{1}^{\tilde{\bm{\pi}}^{*}}$, which is the value of the lower-level policy in T1T-1 periods (and zero terminal value after that).
@@
[…]
@@
    Solve the lower-level MDP with frozen states to obtain a policy $\tilde{\bm{\pi}}^{*}$ and its value $J_{1}^{*}$. Since the lower-level problem is a finite horizon MDP, it can be solved exactly using T1T-1 steps of VI. An alternative way of viewing the lower-level problem is that it is $|\mathcal{X}|$ independent MDPs, which we can potentially solve in parallel. Note that Algorithm 2 emphasizes this by looping over the slow state $x\in\mathcal{X}$ in Line 1.
@@
[…]
@@
2
31exfor $k^{\prime}=1,2,\ldots,k$ do
4
5  0.5em for $s$ in the state space $\mathcal{S}$ do
6
7     0.5em $U^{k^{\prime}}(s)=\max_{a}r(s,a)+\gamma\,\mathbb{E}\bigl[U^{k^{\prime}-1}(f(s,a,w))\bigr].$
8     0.5em
@@
12
130.5em for $s$ in the state space $\mathcal{S}$ do
14
15  0.5em $\nu^{k}(s)=\operatorname*{arg\,max}_{a}r(s,a)+\gamma\,\mathbb{E}\bigl[U^{k}(f(s,a,w))\bigr]$.
16  0.5em
@@
[…]
2
31ex
4for each slow state $x\in\mathcal{X}$ do
5
6  0.5em for $t=T-1,T-2,\ldots,1$ do
7
8     0.5em for each fast state $y\in\mathcal{Y}$ do
9
10        0.5em $J^{*}_{t}(x,y)=\max_{a}r(x,y,a)+\gamma\,\mathbb{E}\bigl[J^{*}_{t+1}(x,f_{\mathcal{Y}}(x,y,a,w))\bigr]$.
11
12        0.5em $\tilde{\pi}_{t}^{*}(x,y)=\operatorname*{arg\,max}_{a}r(x,y,a)+\gamma\,\mathbb{E}\bigl[J^{*}_{t+1}(x,f_{\mathcal{Y}}(x,y,a,w))\bigr]$.
13        0.5em
@@
19
201exfor $k^{\prime}=1,2,\ldots,k$ do
21
22  0.5em for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
23
[…]
25     0.5em
@@
29
300.5em for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
31
[…]
33  0.5em
@@
$\text{Cost}_{\text{FSVI}}(k)=\;\mathcal{O}\!\Bigl(\,\underbrace{|\mathcal{X}|\,|\mathcal{Y}|^{2}\,|\mathcal{A}|\,T\;+\;|\mathcal{S}|^{2}\,T}_{\begin{subarray}{c}\text{one-time lower-level}\\ \text{pre–processing cost}\end{subarray}}\;+\;\underbrace{k\,|\mathcal{S}|^{2}\,|\mathcal{A}|}_{\begin{subarray}{c}\text{upper-level}\\ \text{iterations}\end{subarray}}\,\Bigr).$
@@
$=r(x_{0},y_{0},a)+\gamma\bigl(\tilde{H}^{T-1}\,\bm{0}\bigr)(x_{1},y_{1}),$ (17)
where $\bm{0}$ is the all-zero value function. The difference between (16) and (17) can be interpreted as follows: in the former, we follow a lower-level policy that is aware of a terminal value $U^{*}$ (but exclude that value when defining the $T$-horizon reward), while in the latter, we follow a lower-level policy that sees zero terminal reward at the end of the T1T-1 periods.
@@
[…]
@@ A General Lemma
$\textnormal{reward error (\& end-of-horizon effect)}\;+\;\textnormal{upper-level freeze error}\;+\;\textnormal{$V$-approximation error},$
@@
[…]
@@
The above procedure can be converted into a simple visualization. In Figure 5, we plot regret versus computational cost. This is done for both FSVI for different values of $T$ and also for standard value iteration (Base VI). Each colored curve traces the regret-cost trade‑off generated by varying $k$ for a distinct $T$ (though $k$ is not shown). Imposing the computational budget corresponds to drawing the vertical line $x=B$ and where this line intersects a curve, one can read off the best attainable regret given budget $B$ for the algorithm represented by that curve. By scanning the intersection points and selecting the one with the lowest regret, one can identify the best algorithm to use, whether it is FSVI with a particular $T$ or simply Base VI. Hence, the plot allows us to visually solve the constrained optimization problem of minimizing regret under a given computational budget. This exercise leads us to interesting takeaways illustrated in Figure 5:
@@
Input: State distribution $\rho$, number of iterations $k$, number of state samples NN, number of next-state samples $M$, and a regression procedure Regress.
1
@@
5
6  0.5em Sample a dataset $\mathcal{D}=\{s^{(i)}\}_{i=1}^{N}\sim\rho$
7
8  0.5em for $s^{(i)}\in\mathcal{D}$ do
9
10     0.5em $u^{(i)}=\max_{a}\,r(s^{(i)},a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k^{\prime}-1}\bigl(s^{(i,a,j)}\bigr)$, where $s^{(i,a,j)}$ is sampled from the generative model, conditional on state $s^{(i)}$ and action $a$.
11     0.5em
@@
13
14  0.5em $U^{k^{\prime}}=\texttt{Regress}(\{(s^{(i)},u^{(i)})\}_{i=1}^{N})$
15  0.5em
@@
17
180.5em for $s\in\mathcal{S}$ do
19
20  0.5em $\nu^{k}(s)=\operatorname*{arg\,max}_{a}\,r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k}\bigl(s^{(a,j)}\bigr)$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
21  0.5em
@@
Input: State distribution $\rho$, number of iterations $k$, number of state samples NN, number of next-state samples $M_{l}$ (lower level), $M_{u}$ (upper level), and a regression procedure Regress.
1
[…]
@@
31exInitialize $\tilde{J}_{T}\equiv 0$.
4
50.5em
6for $t=T-1,T-2,\ldots,1$ do
7
8  0.5em Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
9
10  0.5em for each $(x^{(i)},y^{(i)})\in\mathcal{D}$ do
11
12     0.5em $J_{t}^{(i)}=\max_{a\in\mathcal{A}}\,r\bigl(x^{(i)},y^{(i)},a\bigr)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}\tilde{J}_{t+1}\bigl(x^{(i)},y^{(i,a,j)}\bigr)$ where $y^{(i,a,j)}$ is sampled from the generative model, conditional on state $(x^{(i)},y^{(i)})$ and action $a$.
@@
14
15  0.5em $\tilde{J}_{t}=\texttt{Regress}\Bigl(\bigl\{\bigl((x^{(i)},y^{(i)}),J_{t}^{(i)}\bigr)\bigr\}_{i=1}^{N}\Bigr)$
@@
17
181exLet $\tilde{\bm{\pi}}^{*}=(\pi_{1}^{*},\ldots\pi^{*}_{T-1})$, with $\pi_{t}^{*}(s)=\operatorname*{arg\,max}_{a}\,r(s,a)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}\,\tilde{J}_{t+1}\bigl(s^{(a,j)}\bigr)$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
191exInitialize $\hat{V}^{0}$ arbitrarily.
201exfor $k^{\prime}=1,2,\ldots,k$ do
21
22  0.5em Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
23
24  0.5em for each $(x^{(i)},y^{(i)})\in\mathcal{D}$ do
25
26     0.5em ${V}^{(k^{\prime})}=\max_{a\in\mathcal{A}}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}{V}^{k^{\prime}-1}\left(x^{(i,a,j)},y^{(i,a,j)}\right)\right]$ where $x_{T}^{(i,a,j)}$, $y_{T}^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled using a $T$-step transition conditioned on $x^{(i)}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $\tilde{J}_{1}$.
@@
28
29  0.5em $\tilde{V}^{k^{\prime}}=\texttt{Regress}\Bigl(\bigl\{\bigl((x^{(i)},y^{(i)}),{V}^{(i)}\bigr)\bigr\}_{i=1}^{N}\Bigr)$
@@
31
321exfor $s_{0}=(x_{0},y_{0})\in\mathcal{S}$ do
33
34  0.5em $\tilde{\mu}^{k}(x_{0},y_{0})=\operatorname*{arg\,max}_{a\in\mathcal{A}}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}\tilde{V}^{k}\left(x_{T}^{(i,a,j)},y_{T}^{(i,a,j)}\right)\right]$ where $x_{T}^{(i,a,j)}$, $y_{T}^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled from a $T$-step transition as above.
@@
  • {Lin, Deep}-FVI: Fitted value iteration (Munos and Szepesvári, 2008) using either linear function approximator (Lin-FVI) or a three-layer neural network function approximator (Deep-FVI). This algorithm is also known as least-squares Monte Carlo, popularized by Longstaff and Schwartz (2001) and frequently used in the operations literature (see, e.g., Nadarajah et al. (2017)).
@@
We test these algorithms across three domains, summarized in Table 2. Each domain uses a high discount factor of $\gamma=0.995$, corresponding to an effective planning horizon of approximately $1/(1-\gamma)=200$ steps. We describe each experiment below; in each section, we use self-contained notation. Detailed experimental settings are given in Appendix B.